# Equilibrium Portfolio Selection under Utility-Variance Analysis of Log Returns in Incomplete Markets

### v1

@@ 4. $\mathrm{d)}$
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,R_{t}^{\rho},\pi^{t,\varepsilon})-J(t,R_{t}^{\rho},\pi^{\rho})\right)$
@@ 6 Numerical Examples
Here, we choose $\theta(X):=X^{+}\wedge 10000$ to ensure that Assumptions 4.1, 4.2, and 5.2 are satisfied. Similar to dai_dynamic_2021, we fix the parameter values: $r=0.017$, $\sigma=0.15$, $X_{0}=\bar{X}=0.273$, $\lambda=0.27$ and $\nu=0.065$. Moreover, we set $\zeta=1$, $\gamma=0.1$ and $T=2$ for the base model. We discretize the horizon $[0,T]$ into time grid $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{i}=i\Delta t$ for $i=0,1,\dots,N$ and $\Delta t=T/N$. Henceforth, for notational simplicity, we use the subscript $n$ to denote the processes evaluated at time $t_{n}$. The truncated Ornstein–Uhlenbeck process $X$ is approximated by the Euler scheme:
@@
Figure 2 presents a sample trajectory of the learned solution $Y_{t}$, the associated equilibrium strategy $\hat{\pi}_{t}$ and the factor process $X_{t}$. It is observed that the sample path of $Y$ exhibit an overall downward trend over time. In our framework, the process $Y$ satisfies
@@
Intuitively, $\mathbb{E}_{s}\left[U(R_{T}-R_{s})\right]$ decreases as time evolves, leading to a decrease in $Y_{s}$. Therefore, $\mathbb{E}_{s}\left[U^{\prime}(R_{T}-R_{s})\right]$ increases as $Y_{s}$ declines. As the investment horizon approaches, the expected marginal utility rises; consequently, the investor becomes more inclined to invest. This also explains why the equilibrium strategy $\hat{\pi}$ in the mean–variance problem in dai_dynamic_2021 remains constant when $\rho=0$, regardless of the volatility of the factor $X_{t}$, since the expected marginal utility is constant in that case.
@@
Figure 3 plots the mean of the strategy $\hat{\pi}$ using $1000$ trajectories under different values of $\gamma$ and $\zeta$. As illustrated, a higher $\gamma$ corresponds to a smaller investment proportion $\hat{\pi}$, which is intuitively reasonable. This behavior reflects that a larger $\gamma$ indicates the agent places a higher weight on the variance term, leading the investor to favor stability over high returns. Moreover, a higher $\zeta$ also leads to a smaller investment proportion $\hat{\pi}$, which is also intuitively reasonable as $\gamma$ represents the constant risk aversion in the CARA utility function.
@@
We then simulate 1,000 trajectories of the log-return process under each strategy and compute the sample mean and variance of $R_{T}-R_{t}$ at any time $t=n\Delta t$, $n=0,1,2,\dots,N$. It can be observed that, by incorporating the variance term in the $t$-functional, our equilibrium strategy reduces the conditional variance at the cost of a very small decrease in conditional expected utility.
Figure 5: The mean of $\hat{\pi}_{t}$ and $Y_{t}$ under trading constraint with different $\rho$.
@@
Table 1: Comparison of Approximation and Constrained Equilibrium Strategies
Table 1 compares the approximate Nash equilibrium strategy $\hat{\pi}^{\rho}$ with the equilibrium strategy $\hat{\pi}$ under a trading constraint at five discrete time point (t=0.0, 0.5, 1.0, 1.5, 2.0)(t=0.0,\;0.5,\;1.0,\;1.5,\;2.0) with $\rho=-0.31$ and $\gamma=1$ fixed [^7] In this case, since the loss from the variance term is relatively small, we use $\gamma=1$ instead of $0.1$ to make the numerical results more intuitive. . The last column reports the relative errors in per mille (‰). The results show that the discrepancies between the two strategies are negligible, and the relative errors remain extremely small across all time points. Furthermore, as $t\rightarrow T$, the relative error exhibits a slightly decreasing trend. This indicates that the expected payoff generated by the approximate Nash equilibrium strategy closely matches that of the true constrained equilibrium.
@@ Appendix A The Energy Inequality and the John-Nirenberg Inequality
$\mathbb{E}_{\tau}\left[\left(\int_{\tau}^{T}Z^{2}_{s}\mathrm{d}s\right)^{n}\right]\leq n!\|Z\|_{\textup{BMO}}^{2n},\quad a.s..$
@@
If $\|Z\|_{\textup{BMO}}<1$, for every stopping time $\tau$
$\mathbb{E}_{\tau}\left[\exp\left(\int_{\tau}^{T}Z_{s}^{2}\mathrm{d}s\right)\right]\leq\frac{1}{1-\|Z\|_{\textup{BMO}}^{2}},\quad a.s..$
@@ Appendix B Sufficient Conditions for the Existence of Solution to BSDEs
satisfies the condition AB if there exist a deterministic function $l\in L^{1}([0,T])$ and a set of vectors $a_{1},\dots,a_{K}\in\mathbb{R}^{N}$ positively spanning $\mathbb{R}^{N}$, such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2},\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\dots,K.$ (B.2)
@@
$L_{k}:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{d},\quad k=1,\dots,K,$
@@
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2}+a_{k}^{\top}z\,L_{k}(t,x,z),\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\dots,K.$ (B.3)
@@
A set of nonzero vectors $a_{1},\dots,a_{K}$ in $\mathbb{R}^{N}$ (with $K>N$) is said to positively span $\mathbb{R}^{N}$ if, for each $a\in\mathbb{R}^{N}$, there exist nonnegative constants $\lambda_{1},\dots,\lambda_{K}$ such that
@@
$|q^{i}(t,x,y,z)|$ $\leq C_{n}\left(1+\sum_{j=1}^{i}|z^{j}|^{2}\right),$ $i=1,\dots,N,\qquad\textnormal{(quadratic-triangular)}$
@@ Appendix C The Challenge of our BSDE System when $\rho\neq 0$
$\frac{\zeta}{2}z^{2}-\rho\zeta z\cdot\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}+\frac{\zeta+1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}$
[…]
$\frac{1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}.$
The first approach typically requires a continuity condition on $y$ to guarantee a contraction mapping that
[…]
which is not satisfied in our case as $y$ is coupled with the quadratic term of $z,\tilde{z}$.
The second approach relies on the AB condition or the wAB condition (see Appendix B) for the convergence of a subsequence of approximated solutions to the approximated BSDE system. Specifically, there exist a deterministic function $l\in L^{1}[0,T]$ and $a_{1}\cdots a_{k}$ that positively span $\mathbb{R}^{N}$ such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}|a_{k}^{\top}z|^{2}+a_{k}^{\top}z\,L_{k}(t,x,z).$
@@
Here, $L_{i}(t,x,z,\tilde{z})\leq C(1+|z|+|\tilde{z}|),\;i=1,2$ for a constant $C$.
[…]
@@
Substituting it into wAB condition, we obtain
$\frac{\zeta}{2}b^{2}+-\left(\rho^{2}\zeta+\frac{\rho^{2}}{2}\right)b+\frac{\zeta+1}{2}\rho^{2}\leq 0,\quad\left(\frac{\zeta}{2}-\frac{\zeta^{2}\rho^{2}}{2(\zeta+1)}\right)b^{2}-\frac{\rho^{2}\zeta^{2}}{2(\zeta+1)^{2}}b^{3}\leq 0.$ (C.1)
and
@@
It then follows from (C.1) that
$\rho^{4}(\zeta+\frac{1}{2})^{2}-\zeta(\zeta+1)\geq 0,\quad\frac{\zeta+1}{\zeta}\leq\frac{(\zeta+1)^{2}}{\rho\zeta}-(\zeta+1)\leq b\leq\frac{\rho^{2}\zeta+\frac{\rho^{2}}{2}+\sqrt{\rho^{4}(\zeta+\frac{1}{2})^{2}-\zeta(\zeta+1)}}{\zeta}\leq\frac{\zeta+1}{\zeta},$
which result in $\rho=1$ and $b=\frac{\zeta+1}{\zeta}>0$. Substituting $\rho=1$ into (C.2), we get that $c>0$ by (C.1). Consequently, we have
$\frac{\zeta+1}{2}c^{2}-\left(\zeta+\frac{1}{2}\right)c+\frac{\zeta}{2}\geq 0,\quad c\geq\frac{\zeta}{\zeta+1},$
which then imply $c\geq 1$.
As we have shown that, for any $a_{i}=(m,n)\in\{a_{1},\cdots,a_{k}\}$ chosen in wAB condition, $m$ and $n$ must satisfy one of the three conditions

### v2

@@ 4. $\mathrm{d)}$
$\limsup\limits_{\varepsilon\rightarrow 0}\frac{1}{\varepsilon}\left(J(t,\pi^{t,\varepsilon})-J(t,\pi^{\rho})\right)$
@@ 6 Numerical Examples
Here, we choose $\theta(X):=X^{+}\wedge 10000$ to ensure that Assumptions 4.1, 4.2, and 5.2 are satisfied. Similar to dai_dynamic_2021, we fix the parameter values: $r=0.017$, $\sigma=0.15$, $X_{0}=\bar{X}=0.273$, $\lambda=0.27$ and $\nu=0.065$. Moreover, we set $\zeta=1$, $\gamma=0.1$ and $T=2$ for the base model. We discretize the horizon $[0,T]$ into time grid $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{i}=i\Delta t$ for $i=0,1,\cdots,N$ and $\Delta t=T/N$. Henceforth, for notational simplicity, we use the subscript $n$ to denote the processes evaluated at time $t_{n}$. The truncated Ornstein–Uhlenbeck process $X$ is approximated by the Euler scheme:
@@
Figure 2 presents a sample trajectory of the learned solution $Y_{t}$, the associated equilibrium strategy $\hat{\pi}_{t}$ and the factor process $X_{t}$. It is observed that the sample path of $Y$ exhibits an overall downward trend over time. In our framework, the process $Y$ satisfies
@@
Intuitively, $\mathbb{E}_{s}\left[U(R_{T}-R_{s})\right]$ decreases as time evolves, leading to a decrease in $Y_{s}$. Therefore, $\mathbb{E}_{s}\left[U^{\prime}(R_{T}-R_{s})\right]$ increases as $Y_{s}$ declines. As the investment horizon approaches, the expected marginal utility rises; consequently, the investor becomes more inclined to invest. This also explains why the equilibrium strategy $\hat{\pi}$ in the mean–variance problem in dai_dynamic_2021 remains constant when $\rho=0$, regardless of the volatility of the factor $X_{t}$ because the expected marginal utility is constant in that case.
@@
Figure 3 plots the mean of the strategy $\hat{\pi}$ using $1000$ trajectories under different values of $\gamma$ and $\zeta$. As illustrated, a higher $\gamma$ corresponds to a smaller investment proportion $\hat{\pi}$, which is intuitively reasonable. This behavior reflects that a larger $\gamma$ indicates the agent places a higher weight on the variance term, leading the investor to favor stability over high returns. Moreover, a higher $\zeta$ also leads to a smaller investment proportion $\hat{\pi}$, which is also intuitively reasonable as $\gamma$ represents the constant risk aversion in the CARA utility function..
@@
We then simulate 1,000 trajectories of the log-return process under each strategy and compute the sample mean and variance of $R_{T}-R_{t}$ at any time $t=n\Delta t$, $n=0,1,2,\cdots,N$. It can be observed that, by incorporating the variance term in the $t$-functional, our equilibrium strategy reduces the conditional variance at the cost of a very small decrease in conditional expected utility.
Figure 5: The mean of $\hat{\pi}_{t}$ and the factor $X_{t}$ under trading constraint with different $\rho$.
@@
Table 1: Comparison of Approximation and Constrained Equilibrium Strategies (six significant digits)
Table 1 compares the approximate Nash equilibrium strategy $\hat{\pi}^{\rho}$ with the equilibrium strategy $\hat{\pi}$ under a trading constraint at five discrete time point (t=0.0, 0.5, 1.0, 1.5, 2.0)(t=0.0,\;0.5,\;1.0,\;1.5,\;2.0) with $\rho=-0.31$ and $\gamma=1$ fixed [^10] In this case, because the loss from the variance term is relatively small, we use $\gamma=1$ instead of $0.1$ to make the numerical results more intuitive. . The last column reports the relative errors defined as $\left|\frac{\hat{\pi}^{\rho}(t)-\hat{\pi}(t)}{\hat{\pi}(t)}\right|$ in per mille (‰). The results show that the discrepancies between the two strategies are negligible, and the relative errors remain extremely small across all time points. Furthermore, as $t\rightarrow T$, the relative error exhibits a slightly decreasing trend. This indicates that the expected payoff generated by the approximate Nash equilibrium strategy closely matches that of the true constrained equilibrium.
7 Concluding Remarks
This paper investigates a time-inconsistent portfolio optimization problem that integrates expected utility maximization with variance penalization of log returns in an incomplete market. For the general correlated market—where direct analysis becomes intractable—we construct an approximate time-consistent equilibrium. To the best of our knowledge, there are only a few results on time inconsistency in incomplete markets, and we briefly summarize four representative examples.
The work yan2019open investigates the mean–variance problem under stochastic volatility. Another example, mentioned in the introduction, is Gu_2020, which studies the equilibrium strategy for a utility–deviation risk control problem. A third example is liang2025dynamic, where the authors analyze equilibrium strategies under weighted utilities. Yet another related work is mbodji2025portfolio, which examines a portfolio selection problem in a stochastic factor model with non-constant time discounting, focusing on the special case $\rho=1$.
These studies adopt different methodologies tailored to their respective model structures, and the techniques are typically not directly transferable across problems, reflecting the substantial difficulty of analyzing time inconsistency in incomplete markets. Our discussion on the approximate equilibrium aims to provide a more flexible and unified framework, which may offer a new perspective for handling a broader class of time inconsistent control problems in general market models and motivate some relevant future studies.
@@ Appendix A The Energy Inequality and the John-Nirenberg Inequality
$\mathbb{E}_{\tau}\left[\left(\int_{\tau}^{T}Z^{2}_{s}\mathrm{d}s\right)^{n}\right]\leq n!\|Z\|_{\textup{BMO}}^{2n}\quad a.s..$
@@
If $\|Z\|_{\textup{BMO}}<1$, for every stopping time $\tau$,
$\mathbb{E}_{\tau}\left[\exp\left(\int_{\tau}^{T}Z_{s}^{2}\mathrm{d}s\right)\right]\leq\frac{1}{1-\|Z\|_{\textup{BMO}}^{2}}\quad a.s..$
@@ Appendix B Sufficient Conditions for the Existence of Solution to BSDEs
satisfies the condition AB if there exist a deterministic function $l\in L^{1}([0,T])$ and a set of vectors $a_{1},\cdots,a_{K}\in\mathbb{R}^{N}$ positively spanning $\mathbb{R}^{N}$ such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2},\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\cdots,K.$ (B.2)
@@
$L_{k}:[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{N\times d}\to\mathbb{R}^{d},\quad k=1,\cdots,K,$
@@
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}\lvert a_{k}^{\top}z\rvert^{2}+a_{k}^{\top}z\,L_{k}(t,x,z),\qquad\text{for all }(t,x,y,z)\text{ and }k=1,\cdots,K.$ (B.3)
@@
A set of nonzero vectors $a_{1},\cdots,a_{K}$ in $\mathbb{R}^{N}$ (with $K>N$) is said to positively span $\mathbb{R}^{N}$ if, for each $a\in\mathbb{R}^{N}$, there exist nonnegative constants $\lambda_{1},\cdots,\lambda_{K}$ such that
@@
$|q^{i}(t,x,y,z)|$ $\leq C_{n}\left(1+\sum_{j=1}^{i}|z^{j}|^{2}\right),$ $i=1,\cdots,N,\qquad\textnormal{(quadratic-triangular)}$
@@ Appendix C The Challenge of our BSDE System when $\rho\neq 0$
For the coupled BSDE system in (4.2), the drift term $f=(f_{1},f_{2})$ can be computed as
$f_{1}=-\left(\frac{\zeta}{2}|\rho Z+\hat{u}|^{2}+\frac{\zeta(1-\rho^{2})}{2}|Z|^{2}-a(t,\hat{\pi})\right)=h_{1}+g_{1}$
where $h_{1}$ is the quadratic term in $f_{1}$ w.r.t $z$ and $\tilde{z}$
$h_{1}=-\frac{\zeta}{2}z^{2}+\rho\zeta z\cdot\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}-\frac{\zeta+1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}$
and $g_{1}$ collects the linear and constant terms
$g_{1}=$ $-\zeta\rho z\frac{e^{-\zeta y}\theta(t,x)}{(\zeta+1)e^{-\zeta y}+\gamma}+r(t,x)+\theta(t,x)\cdot\frac{e^{-\zeta y}\theta(t,x)-\zeta e^{-\zeta y}\rho z-\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}$
$-\frac{\zeta+1}{2}\left(\frac{e^{-\zeta y}\theta(t,x)}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}+(\zeta+1)\frac{e^{-\zeta y}\theta(t,x)}{(\zeta+1)e^{-\zeta y}+\gamma}\cdot\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}.$
Similarly, $f_{2}=a(t,\hat{\pi})=h_{2}+g_{2}$, where $h_{2}$ is the quadratic term in $f_{2}$ w.r.t. $z$ and $\tilde{z}$
$h_{2}=-\frac{1}{2}\left(\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}.$
and the linear and constant terms are given by
$g_{2}=$ $r(t,x)+\theta(t,x)\cdot\frac{e^{-\zeta y}\theta(t,x)-\zeta e^{-\zeta y}\rho z-\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}$
$-\frac{1}{2}\left(\frac{e^{-\zeta y}\theta(t,x)}{(\zeta+1)e^{-\zeta y}+\gamma}\right)^{2}+\frac{e^{-\zeta y}\theta(t,x)}{(\zeta+1)e^{-\zeta y}+\gamma}\cdot\frac{\zeta e^{-\zeta y}\rho z+\gamma\rho\tilde{z}}{(\zeta+1)e^{-\zeta y}+\gamma}.$
The first approach typically requires a continuity condition in $y$ ensuring that
@@
so that a contraction mapping applies. However, this condition is violated in our setting because $y$ appears inside the quadratic term of $z,\tilde{z}$.
The second approach relies on the AB condition or the wAB condition (see Appendix B) to guarantee the convergence of a subsequence of approximating solutions to the limiting BSDE system. Specifically, assume that the wAB condition holds. Then there exist a deterministic function $l\in L^{1}[0,T]$, vectors $\{a_{1},\cdots,a_{K}\}$ positively spanning $\mathbb{R}^{2}$, and functions $L_{k}:[0,T]\times\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R},\;k=1,\cdots,K$, satisfying $L_{k}\leq C(1+|z|)$ such that
$a_{k}^{\top}f(t,x,y,z)\leq l(t)+\frac{1}{2}|a_{k}^{\top}z|^{2}+a_{k}^{\top}z\,L_{k}(t,x,z),\quad k=1,\cdots,K.$
@@
where $L_{i}(t,x,z,\tilde{z})\leq C(1+|z|+|\tilde{z}|),\;i=1,2$ for a constant $C$.
If b=0(c=0)b=0\;(c=0), the inequality (C.1)((C.2)) fails to hold as $\tilde{z}\;(z)\rightarrow\infty$ and $z\;(\tilde{z})=0$ because $\rho\neq 0$ [^11] Only when $\rho=0$ can vectors of the form (a,0)(-a,0) and (0,d)(0,-d) belong to $\{a_{1},\cdots,a_{k}\}$; see Section 4.2.. We now assume that $b\neq 0$ and $c\neq 0$. Then (C.1) and (C.2) imply
$af_{1}+bf_{2}\leq l(t),\quad\textup{for any}\;(t,x,y,z,\tilde{z})\;\textup{such that}\;az=b\tilde{z}$, (C.3)
and
$cf_{1}-df_{2}\leq l(t),\quad\textup{for any}\;(t,x,y,z,\tilde{z})\;\textup{such that}\;cz=d\tilde{z}$. (C.4)
By substituting $z=\frac{b}{a}\tilde{z}$, the left-hand side of the inequality (C.3) becomes a quadratic polynomial in $\tilde{z}$, while the right-hand side is a deterministic function independent of $z$ and $\tilde{z}$. A necessary condition for the inequality to hold is that the quadratic coefficient of the polynomial is non-positive, namely
$ah_{1}+bh_{2}\leq 0,\quad\textup{for any}\;(t,x,y,z,\tilde{z})\;\textup{such that}\;az=b\tilde{z}$,
Moreover, letting $y\rightarrow\infty$ and $y\rightarrow$-$\infty$ yields
@@
Without loss of generality, set $a=1$. Substituting it into $h_{1}$ and $h_{2}$ with $z=b\tilde{z}$ gives
$\frac{\zeta}{2}b^{2}-\left(\rho^{2}\zeta+\frac{\rho^{2}}{2}\right)b+\frac{\zeta+1}{2}\rho^{2}\leq 0,\quad\left(\frac{\zeta}{2}-\frac{\zeta^{2}\rho^{2}}{2(\zeta+1)}\right)b^{2}-\frac{\rho^{2}\zeta^{2}}{2(\zeta+1)^{2}}b^{3}\leq 0.$ (C.5)
Likewise, from (C.4) we obtain
$ch_{1}-dh_{2}\leq 0,\quad\textup{for any}\;(t,x,y,z,\tilde{z})\;\textup{such that}\;cz=d\tilde{z}$.
Setting $d=1$ and substituting $\tilde{z}=cz$, then letting $y\to\pm\infty$, we get
@@
From (C.5), we deduce
$\left\{\begin{aligned} &\rho^{4}(\zeta+\frac{1}{2})^{2}-\rho^{2}\zeta(\zeta+1)\geq 0,\\ &\frac{\rho^{2}\zeta+\frac{\rho^{2}}{2}-\sqrt{\rho^{4}(\zeta+\frac{1}{2})^{2}-\rho^{2}\zeta(\zeta+1)}}{\zeta}\leq b\leq\frac{\rho^{2}\zeta+\frac{\rho^{2}}{2}+\sqrt{\rho^{4}(\zeta+\frac{1}{2})^{2}-\rho^{2}\zeta(\zeta+1)}}{\zeta}\leq\frac{\zeta+1}{\zeta},\\ &\frac{\zeta+1}{\zeta}\leq\frac{(\zeta+1)^{2}}{\rho^{2}\zeta}-(\zeta+1)\leq b,\end{aligned}\right$.
which result in $\rho=1$ and $b=\frac{\zeta+1}{\zeta}>0$. Substituting $\rho=1$ into (C.6) then yields $c\geq\frac{\zeta}{\zeta+1}$.
We have therefore shown that if $\rho=1$, then any vector $a_{i}=(m,n)\in\{a_{1},\cdots,a_{k}\}$ selected in the wAB condition must satisfy one of the following: